# Which Transformer to Favor: A Comparative Analysis of Efficiency in Vision Transformers

### v3

@@ 1 Introduction
Figure 1: Pareto front (dotted line) of throughput and accuracy. Marker shapes and hues signify the different overarching strategies for making ViT more efficient. Pareto optimal models have a black dot, while the others have a white dot. The size of markers signifies if we fine-tuned using the standard $224$px or the larger $384$px resolution. We mark the baselines ViT-Ti@224 (A), ViT-Ti@384 (B), and ViT-S@224 (C) and also plot the Pareto front for high resolution images (dashed line).
@@
Efforts have been made to reduce the computational load of Transformer models, particularly for resource-constrained settings, such as embedded systems [64, 51]. Researchers have explored numerous strategies, like implementing sparse, local, or kernelized attention mechanisms, as well as token removal criteria to decrease the sequence length. However, selecting the most efficient model that meets certain performance standards remains a challenging task. This task becomes particularly difficult, knowing that “efficiency” can refer to different concepts, like training resources, inference speed, or memory requirements. Moreover, difficulties identifying the most efficient, yet highly performant vision transformers are accentuated by different training and evaluation conditions reported in the literature. It is even more unclear how tradeoffs in efficiency for models proposed for NLP carry over to vision tasks.
@@ 3 Efficient Transformers for Vision
First, we present our methodology for efficiency assessment. Then we briefly describe ViT, as well as a diverse collection of efficient transformers that have been proposed and what these change compared to ViT. These models from the domains of CV and NLP all claim to improve the efficiency of the baseline transformer. We systematically select efficient transformers based on popularity, diversity, and novelty of their approaches. See the supplementary material for more detailes on model selection.
@@ 3.2 Core Elements of ViT
where the softmax is calculated row-wise. This matrix encodes the global interactions between every possible pair of tokens. It is also the reason why the attention mechanism has an inherent computational (space & time) complexity of $\mathcal{O}(N^{2})$. The output sequence of the attention-mechanism is a weighted sum of the input sequence, using the value matrix ($V$):
@@ 3.3 Efficiency-Improving Changes
Low-rank attention leverages the fact, that $Q$, and $K$ in Equation 1 are matrices of shape $N\times d$, which makes ${QK^{\top}\in\mathbb{R}^{N\times N}}$ a matrix of rank $r\leq d\ll N$. The Linformer [65] utilizes this to project the sequence direction of $K$ and $V$ down to dimension $k\ll N$, without reducing the informational content of the attention matrix too much. Similarly, the Nyströmformer [68] uses the Nyström method of matrix decomposition to approximate the matrix $QK^{\top}$. The approximate attention mechanism’s output is then computed with linear complexity by applying the $\operatorname{softmax}$ to each part individually. The approach of XCiT [17] utilizes a transposed attention mechanism:
@@
Fourier attention uses the Fast Fourier Transform (FFT) in the calculation of the attention mechanism to reduce the complexity to $\mathcal{O}(N\log N)$. FNet [30, 48] directly exploits the FFT for $\mathcal{O}(N\log N)$ complexity token mixing, while GFNet [43] utilizes it for a global convolution and AFNO [20] uses an MLP in Fourier space.
@@
Token Removal requires identifying which tokens to remove without losing critical information. To that end, Dynamic ViT [42] uses the Gumbel-softmax for token retention probabilities, and A-ViT [74] learns halting probabilities that weight the token outputs of different depths. In contrast, EViT [34] avoids introducing extra parameters by utilizing the previous layer’s attention matrix.
@@
Summary tokens, condensing the token sequence into a few new tokens, contrasts the approaches so far, which modify the existing tokens. CaiT [59] uses token summary via cross-attention on a single token to gather global information on the classification decision in the last layers. Token Learner [46] creates a set number of summary tokens by employing dynamic sums over the image tokens, and STViT [10] initializes summary tokens with local information by using strided convolutions, then injecting global information into those using cross-attention.
@@
More MLPs can be used to tackle this problem by introducing multiple parameter sets for each MLP block and passing different tokens through blocks with different parameters. Switch Transformer [18] can have multiple times the number of parameters of the ViT without introducing lots of extra computations. In contrast, HiViT [78] exchanges early attention layers for additional MLP blocks.
@@ 4.1 Baselines
To find the most efficient models, we need to quantify their efficiency gains by comparing them to the original ViT [16] and its follow-up versions [57, 58]. We also include ResNet50 [24] in our evaluation as a representative baseline for CNN architectures and a point of reference across papers. All evaluations are based on the ImageNet-1k dataset for classification [15] as it is one of the most commonly used datasets in this domain. Results on other datasets (Stanford Cars [28], Oxford Flowers 102 [39], MIT Places365 [79]) can be found in the supplementary material.
@@ 4.2 Training Pipeline
We compare models on even grounds by training with a standardized pipeline. To reduce bias, our pipeline is relatively simple and only consists of elements commonly used in CV. In particular, we refrain from using knowledge distillation to prevent introducing bias from the choice of teacher model. Any orthogonal techniques, like quantization, sample selection, and others, are not included as they can be applied to every model and would manifest as a systematic offset in the results. To avoid issues from limited training data, we pre-train all models on ImageNet-21k [44].
@@
Pretraining is conducted on ImageNet-21k for 90 epochs at resolutions of $224$ and $192$ pixels, followed by fine-tuning on ImageNet-1k at $224$ and $384$ pixels for 50 epochs. In cases where model training is unstable, we default back to the hyperparameters reported in the corresponding publications. All training is conducted using 4 or 8 NVIDIA A100 GPUs. While most of the models work well with our training pipeline, some did not converge. See the supplementary material for more details.
@@ 4.3 Efficiency Metrics
As we saw, efficiency has multiple dimensions, and therefore we need to capture a variety of metrics. For comparison with other papers and to evaluate their use as proxy metrics for speed and memory requirements, we capture the theoretical metrics of number of parameters and FLOPS. These are used by the community to provide estimates of the model’s representational capacity and computational requirements [41, 55, 43]. However, theoretical metrics do not always correlate with real-world performance, especially when comparing different architectures [14, 4]. Instead, empirical metrics, obtained by running models on hardware, offer more informative evaluations, which is why we track GPU time for training, inference throughput at optimal batch sizes, and VRAM requirements. While these empirical metrics can be sensitive to hardware and software configurations, we ensure consistent evaluations by employing the same setup for all models. We evaluate the variance of our metrics in the supplementary material.
@@ 5.1 Improved Training Pipeline
To validate the fairness of our training pipeline, in Table 1 we compare ImageNet-1k accuracy reported in the original papers (whenever reported) with the one we obtained to see if there are any obvious outliers. Note, that a substantial fraction (13 out of 26) of papers explicitely base their training pipelines on DeiT [57], making them a good fit for training with the updated version in DeiT III [58]. We observe that almost all models trained with this pipeline achieve a higher accuracy; 0.85% on average, with the largest gain being +6.86% for Poly-SA. The models reporting higher performance than what we obtain use knowledge distillation, which we avoid to not induce bias, or a higher resolution, which we find is inefficient.
Most categories contain models that achieve peak accuracy of approximately 85% (see the supplementary material). However, two notable exceptions are observed: the Kernel Attention class, which might pose challenges in optimizing models, and the Fixed Attention class, where the use of a constant attention matrix inherently results in a lower accuracy. Overall, our selection of the training pipeline as a reference for a fair comparison is supported by the improvements in results compared to the original papers, regardless of architectural differences.
Figure 3: Accuracy (red line, right y-axis) and accuracy per parameter (bars, left y-axis) of models ordered by accuracy at a resolution of $224$px. 14 of 47 models of intermediate size are grouped into others. See the supplementary material for the full plot.
@@ 5.2 Number of Parameters
The number of parameters is used in the literature as a proxy metric for tracking the model complexity and the overall computational cost of using a model. When analyzing the parameter efficiency of different models (Figure 3), it is evident that for most smaller models, the accuracy per parameter remains relatively constant at a baseline level of about $4\times 10^{-8}\frac{\%}{\text{param.}}$. This value is approximately halved for larger models, indicating diminishing returns on scaling the model size. For ViT, there is a noticeable drop in accuracy per parameter across the model sizes. While ViT-Ti outperforms models of comparable accuracy, and ViT-S performs on par with similar models, ViT-B slightly underperforms when compared to other, larger models. Significant outliers are the smaller Hybrid Attention models EfficientFormerV2-S0 and CoaT-Ti, which exhibit the highest accuracy per parameter, as well as those Non-attention Shuffling models which incorporte convolutions, like SwiftFormer and FastViT. In the MLP category, we find HiViT to have significantly higher and Switch ViT with significantly lower accuracy per parameter than surrounding models. This suggests that the combination of attention and convolutions allows for the development of highly parameter-efficient models.
@@ 5.3 Speed
Figure 4: Pareto front of finetuning time and accuracy for models which need less than 50 hours for finetuning. We include the full plot in the supplementary material.
@@
We also analyze the Pareto optimality with models finetuned on Stanford Cars [28], Oxford Flowers 102 [39], and MIT Places365 [79] in the supplementary material, observing a high Spearman correlation between the accuracy of different datasets. In particular, the Pareto optimal models of throughput and accuracy are largely the same, which is a strong indicator for the generalizability of our results. Even when assessing throughput on a CPU instead of a GPU, we observe only a high Spearman correlation with only few outliers. Again, we find the Pareto front to be robust to the choice of hardware. See the supplementary material for more information about these results.
@@ 6 Discussion & Conclusion
Figure 6: Suggestions on Which Transformer to Favor in image classification. First, choose the main constraint (1.), speed, or memory, and the phase you want to optimize this constraint at (2.): Inference or training. Based on that, we provide recommendations on which transformer to favor (3.) from the taxonomy classes of Baseline, Sequence Reduction, Hybrid Attention, Non-Attention Shuffling and Sparse Attention. See the supplementary material for a full list of Pareto optimal models for each metric.

### v4

@@ 1 Introduction
Figure 1: Pareto front (dotted line) showing the throughput-accuracy trade-off for efficient Vision Transformers. Markers (varying in shape and hue) represent different efficiency strategies, with black dots highlighting Pareto optimal models. Marker size indicates fine-tuning resolution. Baselines ViT-Ti@224 (A), ViT-Ti@384 (B), and ViT-S@224 (C) are shown, along with the Pareto front for high-resolution images (dashed line).
@@
Efforts have been made to reduce the computational load of Transformer models, particularly for resource-constrained settings, like embedded systems [66, 53]. Researchers have explored numerous strategies, like implementing sparse, local, or kernelized attention mechanisms. However, selecting the most efficient model that meets certain performance standards remains a challenging task, especially since “efficiency” can refer to different concepts, like training resources, inference speed, or memory requirements. Moreover, difficulties identifying the most efficient, yet highly performant vision transformers are accentuated by different training and evaluation conditions reported in the literature. It is even more unclear how tradeoffs in efficiency for models proposed for NLP carry over to vision tasks.
@@ 3 Efficient Transformers for Vision
We present our efficiency assessment methodology and briefly describe ViT, along with various efficient transformer variants. These models from the domains of CV and NLP all claim to improve the efficiency of the baseline transformer. We systematically select efficient transformers based on popularity, diversity, and novelty of their approaches. See Appendix B for more details on model selection.
@@ 3.2 Core Elements of ViT
where the softmax is calculated row-wise. This matrix encodes the global interactions between every possible pair of tokens. It is also the reason why the attention mechanism has an inherent computational (space & time) complexity of $\mathcal{O}(N^{2})$. The output sequence of the attention-mechanism is a weighted sum of the input, using the value matrix ($V$):
@@ 3.3 Efficiency-Improving Changes
Low-rank attention leverages the fact, that ${QK^{\top}\in\mathbb{R}^{N\times N}}$ in Equation 1 is a matrix of rank $r\leq d\ll N$. The Linformer [67] utilizes this to project the sequence direction of $K$ and $V$ down to dimension $k\ll N$, without reducing the informational content of the attention matrix too much. Similarly, the Nyströmformer [71] uses the Nyström method of matrix decomposition to approximate the matrix $QK^{\top}$. The approximate attention mechanism’s output is then computed with linear complexity by applying the $\operatorname{softmax}$ to each part individually. The approach of XCiT [17] utilizes a transposed attention mechanism:
@@
Fourier attention uses the Fast Fourier Transform (FFT) to reduce attention complexity to $\mathcal{O}(N\log N)$. FNet [31, 50] directly exploits the FFT for $\mathcal{O}(N\log N)$ complexity token mixing, while GFNet [45] utilizes it for a global convolution and AFNO [20] uses an MLP in Fourier space.
@@
Token Removal selectively removes tokens while preserving critical information. Dynamic ViT [44] uses Gumbel-softmax for token retention probabilities, and A-ViT [78] learns halting probabilities to weight token outputs at different depths. In contrast, EViT [35] avoids introducing extra parameters by utilizing the previous layer’s attention matrix.
@@
Summary tokens condense the token sequence into a few new tokens, unlike the previous methods that modify existing tokens. CaiT [61] uses cross-attention on a single token to gather global information on the classification decision in the last layers. Token Learner [48] creates a set number of summary tokens by employing dynamic sums over the image tokens, and STViT [10] initializes summary tokens with local information by using strided convolutions, then injecting global information into those using cross-attention.
@@
More MLPs can be used to move compute into these efficient blocks. Switch Transformer [18] introduces multiple parameter sets for different sets of tokens, increasing the number of parameters without significantly introducing extra computations. In contrast, HiViT [83] exchanges early attention layers for additional MLP blocks.
@@ 4.1 Baselines
To find the most efficient models, we need to quantify their efficiency gains by comparing them to the original ViT [16] and its follow-up versions [59, 60]. We also include ResNet50 [24] in our evaluation as a representative baseline for CNN architectures and a point of reference across papers. All evaluations are based on the ImageNet-1k dataset for classification [15] as it is one of the most commonly used datasets in this domain. Results on other datasets (Stanford Cars [29], Oxford Flowers 102 [40], MIT Places365 [84]) can be found in Appendix C.
@@ 4.2 Training Pipeline
We compare models on even grounds by training with a standardized pipeline. To reduce bias, our pipeline is relatively simple and only consists of elements commonly used in CV. In particular, we refrain from using knowledge distillation to prevent introducing bias from the choice of teacher model. Any orthogonal techniques, like quantization, sample selection, and others, are not included as they can be applied to every model and would manifest as a systematic offset in the results. To avoid confounding effects from limited training data, we pre-train all models on ImageNet-21k [46]. While data scarcity impacts performance, we focus on evaluating each architecture’s maximum potential.
@@
Pretraining is conducted on ImageNet-21k for 90 epochs at resolutions of $224$ and $192$ pixels, followed by fine-tuning on ImageNet-1k at $224$ and $384$ pixels for 50 epochs. In cases where model training is unstable, we default back to the hyperparameters reported in the corresponding publications. All training is conducted using 4 or 8 NVIDIA A100 GPUs. While most of the models work well with our training pipeline, some did not converge. See Section A.2 for more details.
@@ 4.3 Efficiency Metrics
As we saw, efficiency has multiple dimensions, and therefore we need to capture a variety of metrics. For comparison with other papers and to evaluate their use as proxy metrics for speed and memory requirements, we capture the theoretical metrics of number of parameters and FLOPS. These are used by the community to provide estimates of the model’s representational capacity and computational requirements [43, 57, 45]. However, theoretical metrics do not always correlate with real-world performance, especially when comparing different architectures [14, 4]. Instead, empirical metrics, obtained by running models on hardware, offer more informative evaluations, which is why we track GPU time for training, inference throughput at optimal batch sizes, and VRAM requirements. While these empirical metrics can be sensitive to hardware and software configurations, we ensure consistent evaluations by employing the same setup for all models. We evaluate the variance of our metrics in Section A.4.
@@ 5.1 Improved Training Pipeline
To validate the fairness of our training pipeline, in Table 1 we compare ImageNet-1k accuracy reported in the original papers (whenever reported) with the one we obtained to see if there are any obvious outliers. Note, that a substantial fraction (13 out of 26) of papers explicitly base their training pipelines on DeiT [59], making them a good fit for training with the updated version in DeiT III [60]. We observe that almost all models trained with this pipeline achieve a higher accuracy; 0.85% on average, with the largest gain being +6.86% for Poly-SA. The models reporting higher performance than what we obtain use knowledge distillation, which we avoid to not induce bias, or a higher resolution, which we find is inefficient.
Most categories contain models that achieve peak accuracy of approximately 85% (see Section D.2). However, two notable exceptions are observed: the Kernel Attention class, which might pose challenges in optimizing models, and the Fixed Attention class, where the use of a constant attention matrix inherently results in a lower accuracy. Overall, our selection of the training pipeline as a reference for a fair comparison is supported by the improvements in results compared to the original papers, regardless of architectural differences.
Figure 3: Accuracy (red line, right y-axis) and accuracy per parameter (bars, left y-axis) of models ordered by accuracy at a resolution of $224$px. 14 of 47 models of intermediate size are grouped into others. See Section D.7 for the full plot.
@@ 5.2 Number of Parameters
The number of parameters is used in the literature as a proxy metric for tracking the model complexity and the overall computational cost of using a model. When analyzing the parameter efficiency of different models (Figure 3), it is evident that for most smaller models, the accuracy per parameter remains relatively constant at a baseline level of about $4\times 10^{-8}\frac{\%}{\text{param.}}$. This value is approximately halved for larger models, indicating diminishing returns on scaling the model size. For ViT, there is a noticeable drop in accuracy per parameter across the model sizes. While ViT-Ti outperforms models of comparable accuracy, and ViT-S performs on par with similar models, ViT-B slightly underperforms when compared to other, larger models. Significant outliers are the smaller Hybrid Attention models EfficientFormerV2-S0 and CoaT-Ti, which exhibit the highest accuracy per parameter, as well as those Non-attention Shuffling models which incorporate convolutions, like SwiftFormer and FastViT. In the MLP category, we find HiViT to have significantly higher and Switch ViT with significantly lower accuracy per parameter than surrounding models. This suggests that the combination of attention and convolutions allows for the development of highly parameter-efficient models.
@@ 5.3 Speed
Figure 4: Pareto front of finetuning time and accuracy for models which need less than 50 hours for finetuning. We include the full plot in Section D.8.
@@
We also analyze the Pareto optimality with models finetuned on Stanford Cars [29], Oxford Flowers 102 [40], and MIT Places365 [84] in Appendix C, observing a high Spearman correlation between the accuracy of different datasets. In particular, the Pareto optimal models of throughput and accuracy are largely the same, which is a strong indicator for the generalizability of our results. Even when assessing throughput on a CPU instead of a GPU, we observe only a high Spearman correlation with only few outliers. Again, we find the Pareto front to be robust to the choice of hardware. See Section D.3 for more information about these results.
@@ 6 Discussion & Conclusion
Figure 6: Suggestions on Which Transformer to Favor in image classification. First, choose the main constraint (1.), speed, or memory, and the phase you want to optimize this constraint at (2.): Inference or training. Based on that, we provide recommendations on which transformer to favor (3.) from the taxonomy classes of Baseline, Sequence Reduction, Hybrid Attention, Non-Attention Shuffling and Sparse Attention. See Section D.10 for a full list of Pareto optimal models for each metric.